# Enhancing Linguistic Competence of Language Models through Pre-training with Language Learning Tasks

### v1

@@ Abstract
Atsuki Yamaguchi, Maggie Mi,  and Nikolaos Aletras School of Computer Science, University of Sheffield, United Kingdom {ayamaguchi1,zmi1,n.aletras}@sheffield.ac.uk
@@ 4.1 Linguistic Competence
Table 2 shows the performance on BLiMP for all models. L2T consistently enhances linguistic competence. In the Disjoint setup, L2T models outperform the Raw baselines at both 500M and 1B scales, achieving overall scores of 80.2 (500M) and 80.8 (1B), respectively, compared to Raw baseline scores of 78.6 and 79.0. Notably, this performance gap widens in the Shared scenario, where L2T models surpass the Raw baselines by 2.8 and 2.3 points. This reinforces our human learning analogy (§2): while Shared Raw mimics “rote learning,” L2T mirrors multi-task acquisition by applying diverse structural lenses to the same data. Our results confirm that linguistic competence depends not just on data volume, but on the diversity of signals applied during training.
@@
We further analyze the effects across different linguistic subfields and phenomena. L2T consistently enhances linguistic competence across model sizes and data scenarios, improving six phenomena across all subfields. Island effects (Island) exhibit the most substantial gains, ranging from 6.9 (1B; Shared) to 11.3 points (1B; Disjoint). Detailed analysis (Appendix D) reveals that nearly all L2T tasks contribute to this gain. This suggests that structured tasks across varying granularity, rather than solely local transformations, provide a structural scaffolding for capturing long-distance dependencies.
Conversely, L2T offers no improvement for determiner-noun agreement (DN Agr) or ellipsis (Ellips). These results serve as a diagnostic of our method: performance for these phenomena appears to reach saturation, as Raw models achieve high accuracy (93+ and 86+, respectively) through implicit learning of frequent patterns Tajeddin and Rahimi (2017); Khullar et al. (2020). This result aligns with the finding of Shah et al. (2024), who report that LMs are substantially accurate on morphological tasks followed by syntactic and semantic tasks. However, limits to this scaffolding remain. Complex phenomena, such as Filler-Gap dependencies (Fill Gap), show no improvement from individual tasks (see Appendix D), indicating that capturing such structures is more challenging and may require targeted discourse-level objectives.
@@ 4.2 General Benchmarks
In the Shared configuration, the impact of L2T varies by model scale. While the 500M model achieves an average improvement of 0.15 points, the 1B model experiences a performance drop of 1.38 points, primarily in CR tasks such as ARC (-4.2). This divergence highlights a tension between linguistic structure learning and factual reinforcement (Fedorenko and Varley, 2016). The Shared Raw baseline encounters the same 42B tokens multiple times, which likely reinforces the retention of factual world knowledge. In contrast, the L2T setup replaces a portion of these repetitions with structured tasks. For the 1B model, the results indicate that the benefits of linguistic stimulation do not fully compensate for the reduced exposure to raw sequences in knowledge-intensive benchmarks. Thus, to perform well on these benchmarks, larger models need to balance structural induction with factual consolidation. [^6] The analysis in Appendix E supports this conclusion by showing that performance on knowledge-intensive benchmarks is sensitive to the ratio of raw text.
@@ Acknowledgements
We would like to thank Anthony Hughes for the valuable feedback. We acknowledge (1) the use of the University of Oxford Advanced Research Computing (ARC) facility: http://dx.doi.org/10.5281/zenodo.22558 and (2) the Isambard-AI National AI Research Resource (AIRR) McIntosh-Smith et al. (2024), which is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. AY is supported by the Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/W524360/1] and the Japan Student Services Organization (JASSO) Student Exchange Support Program (Graduate Scholarship for Degree Seeking Students). MM is supported by the UKRI AI Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].
@@ C.2 Implementation and Training Details
We preprocess datasets with Hugging Face Datasets (v3.2.0) Lhoest et al. (2021). We use PyTorch (v2.3.0) Ansel et al. (2024), FlashAttention-2 (v2.7.3) Dao (2023), and Hugging Face Transformers (v4.49.0) Wolf et al. (2020) for pre-training. For evaluation, we use lm-evaluation-harness Gao et al. (2023) (v0.4.8).
@@ C.3 Evaluation Details: BLiMP Benchmark
Appendix D Effectiveness of Individual Tasks
To examine task-specific contributions to linguistic competence, we pre-train 500M models on 25B task-specific tokens for each of the 14 tasks listed in §2.777Each pre-training dataset originates from the same pool used for the reference data (Raw and L2T). Subsampling ensures that each model processes exactly 25B tokens. Computational constraints limit the total training budget to 25B tokens. However, this allocation is sufficient to cover the window of maximal development Shah et al. (2024), during which models substantially develop cognitive abilities up to approximately 20B tokens. Figure 4 illustrates the BLiMP performance of these models.
Models pre-trained on a single task often maintain competitive performance or demonstrate improvements over the Raw baseline across linguistic subfields. Notably, nine tasks (including Char Count, Last, Masked Word, and Reordering) consistently outperform the Raw model. These tasks likely provide the structural scaffolding that facilitates the acquisition of morphology, syntax, and semantics. Conversely, Masked Char, Space, and One consistently underperform the Raw baseline. The character-level corruption in the Space task shows a 33-point drop in Morphology. We speculate that character-level corruption in tasks like Masked Char and Space might create an unstable training signal when used standalone. The One task, where a model needs to generate a coherent text given a single-word prefix, likely lacks sufficient structured signal for effective learning, although it preserves general capabilities because its objective closely mirrors standard CLM (Figure 6).
@@ Appendix D Effectiveness of Individual Tasks
Finally, this analysis underscores the robustness of the full L2T framework. Despite the inclusion of weaker tasks (namely, Space and One), L2T 100% outperforms most single-task models on BLiMP and general benchmarks (Figure 6). This highlights the benefit of combining diverse tasks; the complementary strengths of the various objectives enhance both linguistic competence and general capabilities. This result mirrors the generalization gains observed when fine-tuning LMs on broad task distributions Wei et al. (2022); Padmakumar et al. (2022); Sanh et al. (2022).
@@
Appendix E Mixing Ratio of Raw and L2T Data
@@
Appendix F Qualitative Analysis
@@ Appendix G License
Appendix H Use of Generative AI Tools

### v2

@@ Abstract
Atsuki Yamaguchi   Maggie Mi   Nikolaos Aletras School of Computer Science, University of Sheffield, United Kingdom {ayamaguchi1,zmi1,n.aletras}@sheffield.ac.uk
@@ 4.1 Linguistic Competence
Table 2 shows the performance on BLiMP for all models. L2T consistently enhances linguistic competence. In the Disjoint setup, L2T models outperform the Raw baselines at both 500M and 1B scales, achieving overall scores of 80.2 (500M) and 80.8 (1B), compared to Raw baseline scores of 78.6 and 79.0. Notably, this performance gap widens in the Shared scenario, where L2T models surpass the Raw baselines by 2.8 and 2.3 points. This reinforces our human learning analogy (§2): while Shared Raw mimics “rote learning,” L2T mirrors multi-task acquisition by applying diverse structural lenses to the same data. Our results confirm that linguistic competence depends not just on data volume, but on the diversity of signals applied during training.
@@
We further analyze the effects across different linguistic subfields and phenomena. L2T consistently enhances linguistic competence across model sizes and data scenarios, improving six phenomena across all subfields. Island effects (Island) exhibit the most substantial gains, ranging from 6.9 (1B; Shared) to 11.3 points (1B; Disjoint). Detailed analysis (§5 and Appendix E) reveals that nearly all L2T tasks contribute to this gain. This suggests that structured tasks across varying granularity, rather than solely local transformations, provide a structural scaffolding for capturing long-distance dependencies.
Conversely, L2T offers no improvement for determiner-noun agreement (DN Agr) or ellipsis (Ellips). These results serve as a diagnostic of our method: performance for these phenomena appears to reach saturation, as Raw models achieve high accuracy (93+ and 86+, respectively) through implicit learning of frequent patterns Tajeddin and Rahimi (2017); Khullar et al. (2020). This result aligns with the finding of Shah et al. (2024), who report that LMs are substantially accurate on morphological tasks followed by syntactic and semantic tasks. However, limits to this scaffolding remain. Complex phenomena, such as Filler-Gap dependencies (Fill Gap), show no improvement from individual tasks (see Appendix E), indicating that capturing such structures is more challenging and may require targeted discourse-level objectives.
@@ 4.2 General Benchmarks
In the Shared configuration, the impact of L2T varies by model scale. While the 500M model achieves an average improvement of 0.15 points, the 1B model experiences a performance drop of 1.38 points, primarily in CR tasks such as ARC (-4.2). This divergence highlights a tension between linguistic structure learning and factual reinforcement (Fedorenko and Varley, 2016). The Shared Raw baseline encounters the same 42B tokens multiple times, which likely reinforces the retention of factual world knowledge. In contrast, the L2T setup replaces a portion of these repetitions with structured tasks. For the 1B model, the results indicate that the benefits of linguistic stimulation do not fully compensate for the reduced exposure to raw sequences in knowledge-intensive benchmarks. Thus, to perform well on these benchmarks, larger models need to balance structural induction with factual consolidation. [^6] The analysis in Appendix F supports this conclusion by showing that performance on knowledge-intensive benchmarks is sensitive to the ratio of raw text.
5 Analysis
Efficacy of Individual Tasks.
To isolate the contributions of specific tasks, we pre-train models on 25B task-specific tokens. As shown in Figure 3, nine tasks (e.g., Char Count, Reordering) consistently outperform the Raw baseline, suggesting they provide critical structural scaffolding for linguistic acquisition. [^7] See Appendix E for the phenomenon-level breakdown. In contrast, character-level tasks such as Space and Masked Char underperform the Raw baseline, with a performance degradation of up to 33 points in Morphology for the Space task. We speculate that these tasks might create an unstable training signal when used standalone. Furthermore, while the One task largely preserves general performance due to the functional overlap of the objective with standard CLM (Figure 6), it provides insufficient structural signal to elicit linguistic gains. Ultimately, the combined L2T framework consistently surpasses the performance of most single-task models on both BLiMP and general benchmarks. This underscores the robustness of the approach and demonstrates that the integration of diverse tasks elicits complementary strengths, a result that parallels the generalization gains observed when fine-tuning LMs on broad task distributions (Wei et al., 2022; Padmakumar et al., 2022; Sanh et al., 2022).
Data Numeric Abilities Conceptual Understanding Fluid Reasoning
Distance Ratio Latent Rep. Zero-Shot RPM
500M Disj. Raw 0.958 0.826 -0.080 0.172 0.312
L2T 0.961 0.830 -0.098 0.184 0.366
Shar. Raw 0.960 0.778 -0.074 0.134 0.292
L2T 0.958 0.847 -0.087 0.201 0.300
1B Disj. Raw 0.947 0.797 -0.128 0.164 0.322
L2T 0.969 0.870 -0.125 0.201 0.352
Shar. Raw 0.970 0.833 -0.125 0.137 0.322
L2T 0.973 0.850 -0.125 0.173 0.298
Table 4: Psychometric evaluation results across cognitive domains. Performance is measured using $R^{2}$ for Numeric Abilities, Spearman $\rho$ for Conceptual Understanding, and Accuracy for Fluid Reasoning. Bold indicates the superior result for a given scale and data regime.
Generalization beyond Linguistic Competence.
We evaluate the impact of L2T on broader cognitive intelligence using the psychometric framework of Shah et al. (2024), covering Numeric Abilities, Conceptual Understanding, and Fluid Reasoning. Table 4 shows that structural stimulation yields consistent gains beyond linguistic competence. For fluid reasoning, L2T substantially enhances Raven Progressive Matrices (RPM) performance (+5.4% for 500M and +3.0% for 1B) in the Disjoint setting. This suggests that L2T enhances the capacity of the model to induce abstract patterns, which is fundamental to fluid intelligence. Furthermore, L2T models consistently demonstrate a stronger Ratio Effect in numeric abilities (e.g., 0.870 for 1B L2T vs. 0.797 for 1B Raw in the Disjoint setting), aligning the representation of magnitude more closely with the processing of humans. Finally, L2T improves conceptual understanding by enhancing zero-shot alignment with the typicality gradients of humans. These results suggest that diverse structured tasks enable the model to develop representations of semantic categorization rather than relying on simple word-level associations.
Discussion.
The primary value of the L2T framework resides in the shift of pre-training from the reproduction of surface patterns to structural induction. As demonstrated in Figure 2, L2T models outperform Raw baselines in linguistic competence as early as 5B tokens, establishing the framework as an efficient mechanism for imparting linguistic inductive biases during the initial stages of pre-training. Unlike instruction tuning, L2T induces structure directly from raw text without external supervision. We expect that the field will adopt the L2T paradigm as a foundational objective to ensure that models grasp the underlying linguistic scaffolding. Nonetheless, a direct comparison with supervised augmentation remains a critical direction. Such an analysis would clarify the advantages of the unsupervised approach of L2T relative to methods that rely on human-annotated data.
@@ 6 Conclusion
Single Seed.
The findings in this study are based on a single training run for each configuration. The high computational cost of pre-training models of 500M and 1B parameters from scratch for 100B tokens made multiple runs prohibitive within the constraints of the available academic infrastructure. However, the consistency of the performance gains across two distinct model scales and two different data regimes (Shared and Disjoint) provides evidence for the robustness of the L2T framework. This indicates that the improvements in linguistic competence result from the systematic impact of structured stimulation rather than sampling variance. Furthermore, the hyperparameters follow the established protocols of Cheng et al. (2024a) to ensure standard training dynamics. To support verification and the scaling of L2T by the research community, the source code and the data transformation pipeline are publicly available at https://github.com/gucci-j/l2t.
Analysis of Individual Tasks.
Assessing the efficacy of individual tasks (§5 and Appendix E) relies on experiments conducted at a reduced scale of 25B tokens. Although this volume of data may not provide a comprehensive profile across the entire pre-training duration, it specifically covers the window of maximal development. During this period, LMs typically demonstrate a linear improvement in cognitive and linguistic abilities (Shah et al., 2024). Because performance gains tend to plateau following this initial phase, this interval represents a theoretically grounded window for the assessment of task impact. This targeted analysis enables the identification of essential structural signals while maintaining computational feasibility.
@@ Acknowledgements
We would like to thank Anthony Hughes for the valuable feedback. We acknowledge (1) the use of the University of Oxford Advanced Research Computing (ARC) facility: http://dx.doi.org/10.5281/zenodo.22558 and (2) the Isambard-AI National AI Research Resource (AIRR) McIntosh-Smith et al. (2025), which is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. AY is supported by the Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/W524360/1] and the Japan Student Services Organization (JASSO) Student Exchange Support Program (Graduate Scholarship for Degree Seeking Students). MM is supported by the UKRI AI Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1]. NA is partly supported by the EPSRC [grant number EP/Y009800/1].
@@ C.2 Implementation and Training Details
We preprocess datasets with Hugging Face Datasets (Lhoest et al., 2021, v3.2.0). We use PyTorch (Ansel et al., 2024, v2.3.0), FlashAttention-2 (Dao, 2024, v2.7.3), and Hugging Face Transformers (Wolf et al., 2020, v4.49.0) for pre-training. For evaluation, we use lm-evaluation-harness (Gao et al., 2023, v0.4.8).
@@ C.3 Evaluation Details: BLiMP Benchmark
Data BoolQ CB MultiRC RTE WiC WSC
500M Disj. Raw 53.1 39.7 48.6 56.0 49.7 38.5
L2T 40.6 31.0 53.3 50.2 47.5 46.2
Shar. Raw 56.5 27.3 53.9 48.7 50.0 41.3
L2T 59.1 31.2 51.1 51.3 47.3 51.9
1B Disj. Raw 55.7 21.3 49.4 52.0 50.5 44.2
L2T 48.4 33.6 51.3 56.0 48.6 45.2
Shar. Raw 40.7 33.8 45.0 48.7 50.9 59.6
L2T 58.0 25.2 49.4 49.5 51.3 42.3
Table 7: Evaluation results on SuperGLUE tasks for models at the 500M and 1B parameter scales. Metrics reported are accuracy (Acc) for all tasks except CB, which uses the F1 score.
Appendix D Analysis on Additional General Benchmarks
We conduct an additional evaluation on SuperGLUE (Wang et al., 2019) to analyze the broader impact of the L2T framework on Natural Language Understanding. While linguistic competence on BLiMP consistently improves (§4.1), results on SuperGLUE exhibit variability across different data scenarios. This variability likely originates from the tension between structural induction and factual reinforcement (Fedorenko and Varley, 2016) as discussed in §4.2. Knowledge-intensive tasks, such as BoolQ and CB, depend on the factual repetition inherent in standard causal language modeling. Consequently, performance decreases in the Disjoint scenario because the exposure to raw text is reduced. Conversely, L2T benefits tasks that require structural resolution, such as MultiRC, by enhancing sensitivity to syntactic dependencies beyond simple word probabilities. The size of the model also modulates these effects. The limited capacity of models with 500M or 1B parameters likely results in high sensitivity to the ratio of raw text necessary for general reasoning. Ultimately, these results indicate that L2T serves as a targeted linguistic force multiplier. The framework achieves higher effectiveness when balanced with sufficient raw text to anchor general factual reasoning.
Appendix E Efficacy of Individual Tasks
@@
Appendix F Mixing Ratio of Raw and L2T Data
@@
Appendix G Qualitative Analysis
@@ Appendix H License
Appendix I Use of Generative AI Tools